# Search for ultralight dark matter in the SuperMAG high-fidelity dataset

### v1

@@ IV.3 Noise estimation
Once the noise PSD for each stationarity period has been computed, the total variance of $\tilde{X}_{k}^{(n)}(f)$ can be determined by combining the noise PSDs from all the stationarity periods which the coherence chunk $k$ overlaps with, that is
@@ IV.4 Bayesian statistical analysis
To turn these likelihoods into posteriors, we require a prior on the couplings. Refs. [50, 51] utilized the Jeffreys prior, which in this context takes the form
@@
Lastly, as in Refs. [50, 51], we apply a 25% degradation factor to our final upper limits
@@ IV.5 Candidate identification and validation
where $F_{\chi^{2}(\nu)}$ is the cumulative distribution function (CDF) of the $\chi^{2}$-distribution with $\nu$ degrees of freedom. We consider there to be a naive DM signal candidate (with 95% global confidence) at a given frequency if $p_{0}$ is below the threshold $p_{\mathrm{crit}}$ defined by
@@
We now wish to evaluate the robustness of these naive signal candidates. In Refs. [50, 51], the identified signal candidates were evaluated by reperforming the analysis on subsets of the data (either a shorter temporal duration or a subset of the stations) and considering whether the signal persisted. Importantly, this was accomplished by comparing the inferred values of the coupling $\varepsilon$ or $g_{a\gamma}$ between the full and subset analyses. The validation procedure used in Refs. [50, 51] therefore did not utilize the full coherence-chunk-to-coherence-chunk phase/directional information contained in the $z_{(i)k}$ variables, but rather reduced all the information to a single variable. The analysis in this work identified many more signal candidates than the $\mathcal{O}(30)$ identified in Refs. [50, 51], and so we require a significantly more powerful validation procedure. In this work, we therefore design a new validation procedure which utilizes the full information contained in the $z_{(i)k}$ variables.
@@
where $N_{f}[T_{\mathrm{coh}}]$ denotes the number of frequencies in the range corresponding to $T_{\mathrm{coh}}$, and $n_{Q_{j}}[f_{p}]$ denotes the index of $f_{p}$ when the frequencies in this range are sorted in ascending order by their $Q_{j}$ values (so that $n_{Q_{j}}=1$ for the lowest value and $n_{Q_{j}}=N_{f}[T_{\mathrm{coh}}]$ for the highest). The $-1$ in the denominator is present so that $f_{p}$ itself is not included in the distribution of $Q_{j}$ used to estimate the $p$-value of $f_{p}$. Eq. (66) is defined to be a conservative estimate of $p_{j}$; that is, it gives an upper bound on the true value of $p_{j}$. For instance, Eq. (66) will give $p_{j}=1/(N_{f}[T_{\mathrm{coh}}]-1)$ for the frequency with the largest $Q_{j}$ value (though its true value of $p_{j}$ may be smaller). [^12] For most coherence times $T_{\mathrm{coh}}$, typically $N_{f}[T_{\mathrm{coh}}]\approx 5.9\times 10^{3}$, so that Eq. (66) always gives an estimate $p_{j}\geq 1.7\times 10^{-5}$. This technique would therefore not have been viable for estimating $p_{0}$, as we need to be able to estimate $p$-values smaller than $p_{\mathrm{crit}}$. Instead, we must rely on a $\chi^{2}$-distribution, as in Eq. (55). [See also footnote 11.]
@@
Note that among the 2931 DPDM candidates and 1186 axion DM candidates we identified, only a single DPDM candidate ($f_{A^{\prime}}=0.9511\,\textrm{Hz}$) and a single axion DM candidate ($f_{a}=0.0303\,\textrm{Hz}$) have $p_{\mathrm{full}}>0.01$. Both of these candidates have weak global significance ($\sigma(p_{0})<2$) and still exhibit tension with the subset analyses ($p_{\mathrm{full}}<0.05$). We therefore deem that our search finds no robust DM candidates. It may however be useful to check these weak candidates in future SNIPE Hunt data. (The axion candidate lies below the frequency range considered in Ref. [52], but may be probed with additional low-frequency data.) Moreover, we also note that none of the signal candidates found in Refs. [50, 51] appear in Tab. 1, and so we reaffirm the claims of those works that none of their candidates were robust.
@@
Figure 3: Constraints on DPDM (left) and axion DM (right) based on our search of the SuperMAG high-fidelity dataset. In solid darker blue, we show our 95%-credible upper limits on $\varepsilon$ and $g_{a\gamma}$, respectively, as a function DM mass (corresponding Compton frequency is shown on the upper axis). These constraints appear as bands due to frequency-to-frequency fluctuations, and also exhibit a number of narrow spikes, which correspond to potential DM signals. We evaluate these candidates in Sec. IV.5 and find that none constitute robust evidence of DM. In solid lighter blue, we also show sliding averages of the limits, which are computed by smoothing with a Gaussian filter on a log-scale [see Eq. (V)]. In various dashed colors, we show existing constraints on DPDM and axion DM (see Sec. V for descriptions). Our constraints are the leading direct probes of both these DM candidates in the mass range $7\times 10^{-17}\,\textrm{eV}\lesssim m_{\mathrm{DM}}\lesssim 4\times 10^{-15}\,\textrm{eV}$, and our DPDM constraint even surpasses the leading astrophysical constraint in a narrow mass range around $m_{A^{\prime}}\approx 2\times 10^{-15}\,\textrm{eV}$.
@@ V Results and conclusion
Having found no robust evidence for a DM signal, we move on to present and discuss the constraints on DPDM and axion DM derived from our analysis. In Fig. 3, we show our resulting constraints on $\varepsilon$ and $g_{a\gamma}$ in darker blue (with the degradation factors in Eqs. (52) and (53) applied). This bound appears as a band due to frequency-bin-to-frequency-bin fluctutations, and also exhibits several narrow peaks, corresponding to the DM signal candidates identified via the procedure in Sec. IV.5. In lighter blue, we show a sliding average of this bound. We compute this average using a Gaussian filter, similar to the one used in Eq. (29) in the noise estimation procedure, but here we perform the smoothing in log-space
@@ Appendix A Vector spherical harmonics
The explicit expressions for the spherical harmonics which are relevant to the DPDM signal [see Eq. (1)] are
@@ B.1 Stationarity timescale
Let us first consider the latter context. The purpose of calculating the noise spectra $S^{a}_{mn}(f)$ is to ultimately determine the covariance matrix $\Sigma_{k}$ for a given coherence chunk. This covariance matrix is computed by summing over all stationarity periods which overlap with the coherence chunk, as in Eqs. (32) and (33). So long as the coherence chunk of interest spans a duration much longer than a year, the difference between summing over all overlapping years vs. all overlapping months/weeks/days is negligible. However, if the coherence chunk is much shorter than a year, then the former will sum over a significantly longer timespan. In particular, if transient events occur in the same calendar year as (but outside of) the coherence chunk, these will affect the noise calculation in the former case but not in the latter case. Since the coherence time in this work can be as short as $T_{\mathrm{coh}}\sim 10^{6}/(1\,\textrm{Hz})\sim 12\,\mathrm{days}$, this makes a year (and likewise any timescale longer than 12 days) a bad candidate for the stationarity timescale. A week or a day are instead more acceptable timescales in this context.
The stationarity timescale plays a much more significant role in the selection of weights. The purpose of the weighting in Eq. (18) is to downweight noisier stations. This is done on the stationarity timescale, so if a transient event of short duration occurs, then the station will be downweighted for the entire stationarity period in which that transient occurs. Moreover, the extent of the downweighting is determined by the DC noise of the station within the stationarity period in question [see Eqs. (19) and (20)]. This means that transient events will be downweighted less when the stationarity timescale is larger because the contribution of the transient to the overall DC noise of the station will be less. For both of these reasons, it is preferable to have a shorter stationarity timescale.
@@
Figure 4: Dependence of noise spectra $S^{a^{\prime}}_{nn}$ (for various years $a^{\prime}$ and time series $X^{(n)}$) in the DPDM analysis on stationarity timescale. We show the spectra as calculated using stationarity timescale (and corresponding window size $\sigma$): calendar year ($\sigma=10^{4}$, blue), calendar month ($\sigma=1000$, orange), week ($\sigma=200$, green), and day ($\sigma=30$, red). [See Eq. (83) for how the spectrum for a year is defined when the stationarity timescale is not a year.] These large window sizes (corresponding to a frequency width of roughly $3\times 10^{-4}\,\mathrm{Hz}$) are used for illustrative purposes, as they lead to spectra which are much smoother than those used in the actual analysis. Note that the lowest noise levels are typically achieved with weekly or daily timescales.
@@ B.2 Filter window
The second analysis parameter which requires validation is the window size $\sigma$ used in the Gaussian filter defined in Eq. (30). This window determines the scale over which the noise spectra are smoothed in the noise estimation procedure. In our analysis, we are searching for signals of width $10^{-6}f_{\mathrm{DM}}$ in frequency space. We identify such signals by comparing them to the local noise. Our smoothing scale should therefore be much larger than the width of a potential signal so that we are not comparing the signal to a noise estimated from the signal itself. Moreover, $\sigma$ should be large enough that our smoothing procedure averages over enough frequency bins to get an accurate estimate of the noise. On the other hand, if we smooth over a window signficantly larger than the width of the actual features in the noise spectrum, we will end up comparing potential signals to noise levels which do not accurately characterize the noise in their local frequency range. This will manifest as improper statistics of the $z_{(i)k}$ variables, which according to our analysis should be Gaussian with mean zero and $\langle|z_{(i)k}|^{2}\rangle=1$.
@@
Fig. 5 demonstrates that a larger window size leads to poorly behaved statistics for the analysis variables. In particular, the four-point statistics in the $\sigma=60$ case differ from their expected value by as much as $40\%$. Many of the large deviations in Fig. 5 are only of width $\sigma_{\log}$, which is indicative of an anomaly that may only span a few frequency bins (but is broadened by the Gaussian filter). As alluded to earlier, the smaller $\sigma=15$ window size models the noise near these anomalies more locally, thereby leading to four-point statistics which are closer to expectation. The two-point statistics are similar in both cases at higher frequencies, but at lower frequencies the larger window size leads to larger deviations in the two-point statistics. This is because the noise exhibits $1/f^{\alpha}$-behavior (see Fig. 4), so the noise changes much more rapidly at low frequencies, and therefore a smaller window size is more appropriate. The $\sigma=15$ case exhibits $\mathcal{O}(5\%)$ deviations in the two-point statistics and even smaller deviations in the four-point statistics, which indicates that our noise modeling is sufficiently accurate with this choice of window size.

### v2

@@ IV.3 Noise estimation
Once the noise PSD for each stationarity period has been computed, the total variance of $\tilde{X}_{k}^{(n)}(f)$ can be determined by combining the noise PSDs from all the stationarity periods that the coherence chunk $k$ overlaps with, that is
@@ IV.4 Bayesian statistical analysis
To turn these likelihoods into posteriors, we require a prior on the couplings. Refs. [53, 54] utilized the Jeffreys prior, [^13] For a discussion on the choice of different priors in ultralight DM searches, we point the interested reader to Appendix I A of Ref. [4], which compares a Jeffreys prior to a uniform prior. They find the choice of prior only results in an $\mathcal{O}(1)$ rescaling of the constraint. Moreover, they point out that the Jeffreys prior is conceptually preferable due to its reparametrization invariance. which in this context takes the form
@@
Last, as in Refs. [53, 54], we apply a 25% degradation factor to our final upper limits
@@ IV.5 Candidate identification and validation
where $F_{\chi^{2}(\nu)}$ is the cumulative distribution function of the $\chi^{2}$-distribution with $\nu$ degrees of freedom. We consider there to be a naive DM signal candidate (with 95% global confidence) at a given frequency if $p_{0}$ is below the threshold $p_{\mathrm{crit}}$ defined by
@@
We now wish to evaluate the robustness of these naive signal candidates. In Refs. [53, 54], the identified signal candidates were evaluated by reperforming the analysis on subsets of the data (either a shorter temporal duration or a subset of the stations) and considering whether the signal persisted. Importantly, this was accomplished by comparing the inferred values of the coupling $\varepsilon$ or $g_{a\gamma}$ between the full and subset analyses. The validation procedure used in Refs. [53, 54] therefore did not utilize the full coherence-chunk-to-coherence-chunk phase/directional information contained in the $z_{(i)k}$ variables, but rather reduced all the information to a single variable. The analysis in this work identified many more signal candidates than the $\mathcal{O}(30)$ identified in Refs. [53, 54], and so we require a significantly more powerful validation procedure. In this work, we therefore design a new validation procedure that utilizes the full information contained in the $z_{(i)k}$ variables.
@@
where $N_{f}[T_{\mathrm{coh}}]$ denotes the number of frequencies in the range corresponding to $T_{\mathrm{coh}}$, and $n_{Q_{j}}[f_{p}]$ denotes the index of $f_{p}$ when the frequencies in this range are sorted in ascending order by their $Q_{j}$ values (so that $n_{Q_{j}}=1$ for the lowest value and $n_{Q_{j}}=N_{f}[T_{\mathrm{coh}}]$ for the highest). The $-1$ in the denominator is present so that $f_{p}$ itself is not included in the distribution of $Q_{j}$ used to estimate the $p$-value of $f_{p}$. Eq. (66) is defined to be a conservative estimate of $p_{j}$; that is, it gives an upper bound on the true value of $p_{j}$. For instance, Eq. (66) will give $p_{j}=1/(N_{f}[T_{\mathrm{coh}}]-1)$ for the frequency with the largest $Q_{j}$ value (though its true value of $p_{j}$ may be smaller). [^15] For most coherence times $T_{\mathrm{coh}}$, typically $N_{f}[T_{\mathrm{coh}}]\approx 5.9\times 10^{3}$, so that Eq. (66) always gives an estimate $p_{j}\geq 1.7\times 10^{-5}$. This technique would therefore not have been viable for estimating $p_{0}$, as we need to be able to estimate $p$-values smaller than $p_{\mathrm{crit}}$. Instead, we must rely on a $\chi^{2}$-distribution, as in Eq. (55). [See also footnote 14.]
@@
Note that among the 2931 DPDM candidates and 1186 axion DM candidates we identified, only a single DPDM candidate ($f_{A^{\prime}}=0.9511\,\textrm{Hz}$) and a single axion DM candidate ($f_{a}=0.0303\,\textrm{Hz}$) have $p_{\mathrm{full}}>0.01$. Both of these candidates have weak global significance ($\sigma(p_{0})<2$) and still exhibit tension with the subset analyses ($p_{\mathrm{full}}<0.05$). We therefore deem that our search finds no robust DM candidates. It may, however, be useful to check these weak candidates in future SNIPE Hunt data. (The axion candidate lies below the frequency range considered in Ref. [55], but may be probed with additional low-frequency data.) Moreover, we also note that none of the signal candidates found in Refs. [53, 54] appear in Tab. 1, and so we reaffirm the claims of those works that none of their candidates were robust.
@@
Figure 3: Constraints on DPDM (left) and axion DM (right) based on our search of the SuperMAG high-fidelity dataset. In solid darker blue, we show our 95%-credible upper limits on $\varepsilon$ and $g_{a\gamma}$, respectively, as a function of DM mass (corresponding Compton frequency is shown on the upper axis). These constraints appear as bands due to frequency-to-frequency fluctuations, and also exhibit a number of narrow spikes, which correspond to potential DM signals. We evaluate these candidates in Sec. IV.5 and find that none constitute robust evidence of DM. In solid lighter blue, we also show sliding averages of the limits, which are computed by smoothing with a Gaussian filter on a log scale [see Eq. (V)]. In various dashed colors, we show existing constraints on DPDM and axion DM (see Sec. V for descriptions). Our constraints are the leading direct probes of both these DM candidates in the mass range $7\times 10^{-17}\,\textrm{eV}\lesssim m_{\mathrm{DM}}\lesssim 4\times 10^{-15}\,\textrm{eV}$, and our DPDM constraint even surpasses the leading astrophysical constraint in a narrow mass range around $m_{A^{\prime}}\approx 2\times 10^{-15}\,\textrm{eV}$.
@@ V Results and conclusion
Having found no robust evidence for a DM signal, we move on to present and discuss the constraints on DPDM and axion DM derived from our analysis. In Fig. 3, we show our resulting constraints on $\varepsilon$ and $g_{a\gamma}$ in darker blue (with the degradation factors in Eqs. (52) and (53) applied). This bound appears as a band due to frequency-bin-to-frequency-bin fluctutations, and also exhibits several narrow peaks, corresponding to the DM signal candidates identified via the procedure in Sec. IV.5. In lighter blue, we show a sliding average of this bound. We compute this average using a Gaussian filter, similar to the one used in Eq. (29) in the noise estimation procedure, but here we perform the smoothing in log space
@@ Appendix A Vector spherical harmonics
The explicit expressions for the spherical harmonics that are relevant to the DPDM signal [see Eq. (1)] are
@@ B.1 Stationarity timescale
Let us first consider the latter context. The purpose of calculating the noise spectra $S^{a}_{mn}(f)$ is to ultimately determine the covariance matrix $\Sigma_{k}$ for a given coherence chunk. This covariance matrix is computed by summing over all stationarity periods that overlap with the coherence chunk, as in Eqs. (32) and (33). So long as the coherence chunk of interest spans a duration much longer than a year, the difference between summing over all overlapping years vs. all overlapping months/weeks/days is negligible. However, if the coherence chunk is much shorter than a year, then the former will sum over a significantly longer timespan. In particular, if transient events occur in the same calendar year as (but outside of) the coherence chunk, these will affect the noise calculation in the former case but not in the latter case. Since the coherence time in this work can be as short as $T_{\mathrm{coh}}\sim 10^{6}/(1\,\textrm{Hz})\sim 12\,\mathrm{days}$, this makes a year (and likewise any timescale longer than 12 days) a bad candidate for the stationarity timescale. A week or a day are instead more acceptable timescales in this context.
The stationarity timescale plays a much more significant role in the selection of weights. The purpose of the weighting in Eq. (18) is to downweight noisier stations. This is done on the stationarity timescale, so if a transient event of short duration occurs, then the station will be downweighted for the entire stationarity period in which that transient occurs. Moreover, the extent of the downweighting is determined by the white noise level of the station within the stationarity period in question [see Eqs. (19) and (20)]. This means that transient events will be downweighted less when the stationarity timescale is larger because the contribution of the transient to the overall white noise level of the station will be less. For both of these reasons, it is preferable to have a shorter stationarity timescale.
@@
Figure 4: Dependence of noise spectra $S^{a^{\prime}}_{nn}$ (for various years $a^{\prime}$ and time series $X^{(n)}$) in the DPDM analysis on stationarity timescale. We show the spectra as calculated using stationarity timescale (and corresponding window size $\sigma$): calendar year ($\sigma=10^{4}$, blue), calendar month ($\sigma=1000$, orange), week ($\sigma=200$, green), and day ($\sigma=30$, red). [See Eq. (83) for how the spectrum for a year is defined when the stationarity timescale is not a year.] These large window sizes (corresponding to a frequency width of roughly $3\times 10^{-4}\,\mathrm{Hz}$) are used for illustrative purposes, as they lead to spectra that are much smoother than those used in the actual analysis. Note that the lowest noise levels are typically achieved with weekly or daily timescales.
@@ B.2 Filter window
The second analysis parameter that requires validation is the window size $\sigma$ used in the Gaussian filter defined in Eq. (30). This window determines the scale over which the noise spectra are smoothed in the noise estimation procedure. In our analysis, we are searching for signals of width $10^{-6}f_{\mathrm{DM}}$ in frequency space. We identify such signals by comparing them to the local noise. Our smoothing scale should therefore be much larger than the width of a potential signal so that we are not comparing the signal to a noise estimated from the signal itself. Moreover, $\sigma$ should be large enough that our smoothing procedure averages over enough frequency bins to get an accurate estimate of the noise. On the other hand, if we smooth over a window significantly larger than the width of the actual features in the noise spectrum, we will end up comparing potential signals to noise levels that do not accurately characterize the noise in their local frequency range. This will manifest as improper statistics of the $z_{(i)k}$ variables, which according to our analysis should be Gaussian with mean zero and $\langle|z_{(i)k}|^{2}\rangle=1$.
@@
Fig. 5 demonstrates that a larger window size leads to poorly behaved statistics for the analysis variables. In particular, the four-point statistics in the $\sigma=60$ case differ from their expected value by as much as $40\%$. Many of the large deviations in Fig. 5 are only of width $\sigma_{\log}$, which is indicative of an anomaly that may only span a few frequency bins (but is broadened by the Gaussian filter). As alluded to earlier, the smaller $\sigma=15$ window size models the noise near these anomalies more locally, thereby leading to four-point statistics that are closer to expectation. The two-point statistics are similar in both cases at higher frequencies, but at lower frequencies the larger window size leads to larger deviations in the two-point statistics. This is because the noise exhibits $1/f^{\alpha}$-behavior (see Fig. 4), so the noise changes much more rapidly at low frequencies, and therefore a smaller window size is more appropriate. The $\sigma=15$ case exhibits $\mathcal{O}(5\%)$ deviations in the two-point statistics and even smaller deviations in the four-point statistics, which indicates that our noise modeling is sufficiently accurate with this choice of window size.